OpenStudy (anonymous):

How would I find the value of x? 2√X=10?

OpenStudy (welshfella):

first divide both sides by 2 then square both sides to give x = ?

OpenStudy (anonymous):

5?

OpenStudy (jdoe0001):

\(\bf 2\sqrt{x}=10\implies \sqrt{x}=\cfrac{10}{2}\implies x=\left( \cfrac{10}{2}\right)^2\)

OpenStudy (jdoe0001):

divide by 2 and then square both sides to get the radicand out

OpenStudy (anonymous):

20?

OpenStudy (jdoe0001):

20? hmmmm what did you get from the fraction?

OpenStudy (anonymous):

5 and then square it?

OpenStudy (anonymous):

25?

OpenStudy (jdoe0001):

yeap

OpenStudy (jdoe0001):

\(\bf 2\sqrt{x}=10\implies \sqrt{x}=\cfrac{\cancel{10}}{\cancel{2}}\implies x=\left( 5\right)^2\implies x=25\)